OpenStudy (anonymous):

Find the value of Q3 for the following data set. 4 6 8 8 8 9 13 15

OpenStudy (amistre64):

q3 = p75

OpenStudy (amistre64):

there are 8 data points; 75% of 8 = 6; it is that value between the 6th and 7th data points

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quartile

OpenStudy (amistre64):

(9+13)/2 = 22/2 = 11

OpenStudy (anonymous):

oh so i just take the 6th and 7th add them then divide by 2? is tht how it is always?

OpenStudy (amistre64):

that is how this one is since the percentage is an integer; if it had been a decimal, you simply round it up to the next spot

OpenStudy (amistre64):

say it had been 6.3678; then you pick the 7th point

OpenStudy (anonymous):

thank u amistre64

OpenStudy (amistre64):

yep